# Ergodic Risk Measures: Towards a Risk-Aware Foundation for Continual Reinforcement Learning

### v1

@@ Abstract
Continual reinforcement learning (continual RL) seeks to formalize the notions of lifelong learning and endless adaptation in RL. In particular, the aim of continual RL is to develop RL agents that can maintain a careful balance between retaining useful information and adapting to new situations. To date, continual RL has been explored almost exclusively through the lens of risk-neutral decision-making, in which the agent aims to optimize the expected (or mean) long-run performance. In this work, we present the first formal theoretical treatment of continual RL through the lens of risk-aware decision-making, in which the agent aims to optimize a reward-based measure of long-run performance beyond the mean. In particular, we show that the classical theory of risk measures, widely used as a theoretical foundation in non-continual risk-aware RL, is, in its current form, incompatible with the continual setting. Then, building on this insight, we extend risk measure theory into the continual setting by introducing a new class of ergodic risk measures that are compatible with continual learning. Finally, we provide a case study of risk-aware continual learning, along with empirical results, which show the intuitive appeal and theoretical soundness of ergodic risk measures.
@@ 1 Introduction
To date, advances in continual RL have been developed almost exclusively under a risk-neutral paradigm, such that the agent is designed to optimize the expected (or average) long-run performance. Yet, the mere notion of lifelong learning implies the notion of survival, and hence, risk-awareness. That is, the agent needs to first survive indefinitely if it wants to continue learning indefinitely. To this end, we argue that, in real-world settings, if the agent wants to survive, it needs to learn how to act in a risk-aware manner, such that it learns to avoid catastrophic scenarios. In particular, we argue that if an agent that is deployed in the real-world cannot learn to avoid catastrophic scenarios, then it is unlikely that the agent will be allowed to, or perhaps may not even be able to, continue operating indefinitely. Moreover, although an agent may learn to avoid catastrophic scenarios as part of its effort to optimize the expected long-run performance, this is far from guaranteed; there always exists the possibility that the agent may choose to engage in such scenarios, if doing so allows it to optimize the expected long-run performance. This hence motivates the need for risk-aware continual RL agents, who can explicitly learn to prioritize avoiding catastrophic scenarios, even if it comes at the cost of optimizing the expected long-run performance.
In this work, we take the first steps towards developing a risk-aware foundation for continual RL. In particular, we first examine the classical theory of risk measures (e.g. see Chapter 6 of Shapiro et al. [2009]), which has served as a crucial theoretical foundation for risk-aware decision-making in non-continual RL, and show that, in its current form, it is inconsistent with the unique demands of continual RL, particularly the stability-plasticity dilemma. Then, building on this insight, we extend risk measure theory into the continual setting by introducing a new class of risk measures, called ergodic risk measures, which are designed to be compatible with continual learning. Finally, using the well-known average-reward Markov decision process (MDP) formulation [Puterman, 1994] as a basis, we provide a case study, along with numerical results, which show the intuitive appeal and theoretical soundness of ergodic risk measures in a continual learning setting. Altogether, these contributions provide, to the best of our knowledge, the first formal theoretical treatment of risk-aware decision-making in a continual (i.e., lifelong) learning setting.
@@ 2.1 Continual Reinforcement Learning
The notions of lifelong learning and endless adaptation in the context of RL have long been studied under different names and perspectives. In recent years, several works (e.g. Khetarpal et al. [2022], Abel et al. [2023], Kumar et al. [2025]) have attempted to unify and frame these diverse sets of works as instances of continual RL. Some of the more common types of RL-related works that can be interpreted as being instances of continual RL include the study of the loss of plasticity in deep RL agents (e.g. Abbas et al. [2023], Dohare et al. [2024]), transfer learning (e.g. Abel et al. [2018], Gimelfarb et al. [2021]), and decision-making in non-stationary environments (e.g. Dick et al. [2014], Luketina et al. [2022]). The term ‘continual RL’ itself was first introduced in Ring [1994], and since then, there have been various works that have looked at extending various aspects of RL into the continual setting, both via the discounted and average-reward MDP formulations.
@@ 2.2 Risk-Aware Reinforcement Learning
The notion of risk-aware learning and decision-making in the context of RL has been studied under various theoretical frameworks, from the well-established expected utility framework [Howard and Matheson, 1972], to the more contemporary framework of risk measures (e.g. Chapter 6 of Shapiro et al. [2009]). In this work, we focus on the latter framework, which originated in the finance literature (e.g. Rockafellar and Uryasev [2000]), but has since been widely integrated into RL-based works (e.g. Bäuerle and Ott [2011]). Of particular importance to the framework of risk measures are the concepts of interpretability, ‘coherence’ [Artzner et al., 1999], and ‘time consistency’ [Boda and Filar, 2006], where the latter two concepts are used to define sub-classes of risk measures that satisfy key mathematical properties which can be meaningful in risk-based decision-making contexts. Traditionally, non-continual risk-aware RL works have aimed to optimize either a ‘static’ (interpretable, coherent, time-inconsistent) risk measure (e.g. Mead et al. [2025]), or a ‘nested (dynamic)’ (hard-to-interpret, coherent, time-consistent) risk measure (e.g. Ruszczyński [2010]). To the best of our knowledge, our work is the first to propose an extension of risk measure theory into the continual learning setting. The case study presented in this work primarily focuses on the conditional value-at-risk (CVaR) risk measure [Rockafellar and Uryasev, 2000], which has been studied extensively in the discounted setting (e.g. Bäuerle and Ott [2011], Mead et al. [2025]), and, to a lesser extent, in the average-reward setting (e.g. Xia et al. [2023], Rojas and Lee [2025]).
@@ 3.1 Continual Reinforcement Learning
Consider a finite MDP, $\mathcal{M}\doteq\langle\mathcal{S},\mathcal{A},\mathcal{R},p\rangle$, where $\mathcal{S}$ is a finite set of states, $\mathcal{A}$ is a finite set of actions, $\mathcal{R}\subset\mathbb{R}$ is a bounded set of rewards, and $p:\mathcal{S}\,\times\,\mathcal{A}\,\times\,\mathcal{R}\,\times\,\mathcal{S}\rightarrow{}[0,1]$ is a probabilistic transition function that describes the dynamics of the environment, such that at each discrete time step, $t=0,1,2,\ldots$, an agent chooses an action, $A_{t}\in\mathcal{A}$, based on its current state, $S_{t}\in\mathcal{S}$, and receives a reward, $R_{t+1}\in\mathcal{R}$, while transitioning to a (potentially) new state, $S_{t+1}$, such that $p(s^{\prime},r\mid s,a)=\mathbb{P}(S_{t+1}=s^{\prime},R_{t+1}=r\mid S_{t}=s,A_{t}=a)$.
A continual RL problem can be viewed as an infinite sequence of MDPs, $\{\mathcal{M}_{k}\}_{k=1}^{\infty}$, such that $\mathcal{M}_{k}\doteq\langle\mathcal{S}_{k},\mathcal{A}_{k},\mathcal{R}_{k},p_{k}\rangle$, where each $\mathcal{M}_{k}$ may differ in its state-space, action-space, reward function, and/or transition dynamics based on some indexing function, $\omega:\mathbb{N}\to\mathbb{N}$, such that $\omega(t)=i$ indicates that at time step $t$, the agent interacts with environment $\mathcal{M}_{i}$. We note that the function $\omega$ need not be known to the agent.
At the heart of continual RL is the stability-plasticity dilemma, which requires the agent to balance two competing demands: retaining useful information learned in prior MDPs to use in later MDPs, and adapting to new streams of experience generated by the differences between the various MDPs. More formally, the agent’s goal in the continual setting is to construct a sequence of stationary policies, $\{\pi_{k}\}_{k=1}^{\infty}$, that optimizes some measure of long-run performance. In this work, we focus on one such measure of long-run performance known as the long-run (or limiting) average-reward, $\bar{r}$, which is defined as follows for a given stationary policy being followed at time $t$, $\pi_{t}$:
$\bar{r}_{\pi_{t}}(s)\doteq\lim_{n\rightarrow{}\infty}\frac{1}{n}\sum_{t=t_{0}}^{n}\mathbb{E}[R_{t}\mid S_{t_{0}}=s,A_{t_{0}:t-1}\sim\pi_{t}],$ (1)
where $t_{0}=\min\{y:\pi_{\omega(y)}=\pi_{t}\}$. MDPs that aim to optimize the long-run average-reward objective (1) are known as average-reward (or average-cost) MDPs [Puterman, 1994]. Our choice of the average-reward MDP formulation in this work (rather than the discounted MDP formulation) follows prior work, such as Sharma et al. [2022] and Kumar et al. [2025], which argue that the average-reward formulation’s emphasis on long-term performance is a natural fit for continual learning settings.
When working with average-reward MDPs, it is common to simplify the expression for the average-reward objective (1) into a more workable form by making certain ergodicity-like assumptions about the Markov chain induced when following policy $\pi_{t}$. To this end, a unichain assumption is typically used when doing prediction (learning) because it ensures the existence of a unique limiting distribution of states, $\mu_{\pi_{t}}(s)\doteq\lim_{t\rightarrow{}\infty}\mathbb{P}(S_{t}=s\mid A_{t_{0}:t-1}\sim\pi_{t})$, that is independent of initial conditions. Similarly, a communicating assumption is typically used for control (optimization) because it ensures the existence of a unique optimal average-reward, $\bar{r}*$, that is independent of initial conditions. Importantly, these ergodicity-like assumptions enable the agent’s objective to be expressed as a stable measure of long-term performance that eventually becomes independent of prior conditions.
@@ 3.2 Risk Measures
Let $(\Omega,\mathcal{F},\mathbb{P})$ denote a probability space, and let $\mathcal{X}$ denote a space of random variables of the form $X:\Omega\to\mathbb{R}$. A risk measure (e.g. Chapter 6 of Shapiro et al. [2009]) is a functional, $\rho:\mathcal{X}\to\mathbb{R}$, that assigns to each random variable, $X\in\mathcal{X}$, a real value representing the degree of risk associated with $X$. In other words, a risk measure is a mapping that quantifies the risk associated with a random variable. The precise interpretation of what ‘risk’ means depends on the risk measure used; different risk measures capture different aspects of the variability and/or tail behavior of a random variable.
@@
In essence, one can think of a risk measure as any functional that captures distributional characteristics of a random variable, typically beyond just its mean (e.g. variance). However, an emphasis is usually placed on deriving risk measures that satisfy certain mathematical properties that can be meaningful in risk-based decision-making contexts. To this end, we now provide informal definitions of the various classes of risk measures used in the context of RL, where each (non-mutually-exclusive) class of risk measures can be thought of as satisfying a specific set of mathematical properties. We provide formal definitions of these risk measures in Appendix A.
Coherent Risk Measures (Definition A.1): A risk measure, $\rho$, is called coherent if it satisfies the following four axioms for all $X,X^{\prime}\in\mathcal{X}$:
  1. 1.
  2. 2.
  3. 3.
  4. 4.
    Subadditivity: $\rho(X+X^{\prime})\leq\rho(X)+\rho(X^{\prime})$.
Static Risk Measures (Definition A.2): A static risk measure evaluates risk at a fixed point in time, without taking into consideration the temporal evolution of information. In RL-based contexts, static risk measures can be useful for quantifying the risk associated with the return at the end of an episode. One of the primary appeals of static risk measures is their interpretability.
Conditional Risk Measures (Definition A.3): A conditional risk measure at time $t$, $\rho_{t}(X)$, is a mapping that evaluates the risk of future outcomes (e.g. at time $N>t$) based on the information available up to and including time $t$.
Dynamic Risk Measures (Definition A.4): A dynamic risk measure is a sequence of conditional risk measures, $\{\rho_{t}(X)\}_{t=0}^{N}$, that allows risk to be tracked and updated as new information becomes available. In RL-based contexts, dynamic risk measures can be useful for capturing the sequential nature of decision-making (i.e., that actions taken at each time step can potentially influence future outcomes, and hence, future risk evaluations).
Time-Consistent Risk Measures (Definition A.5): A dynamic risk measure, $\{\rho_{t}(X)\}_{t=0}^{N}$, is said to be time-consistent if, for all $X,X^{\prime}\in\mathcal{X}$ and all $t<N$,
$\rho_{t+1}(X)\leq\rho_{t+1}(X^{\prime})\implies\rho_{t}(X)\leq\rho_{t}(X^{\prime}).$
Time-consistent risk measures can be appealing because they ensure that if one future outcome is deemed less risky than another at some time step, $t$, then that same outcome is not deemed more risky than the other at any other time step. In RL-based contexts, time-consistent risk measures can be useful because they can induce Bellman-like recursions with appealing dynamic programming-like properties. One way to think about time consistency in RL-based contexts is to ask the question: can the agent change its mind about how risky something is based on new information? If the answer is yes, then there exists a lack of time consistency.
Nested Risk Measures (Definition A.6): A nested risk measure is a dynamic risk measure that is constructed recursively from one-step conditional risk measures:
$\rho_{0:N}(X)\doteq\rho_{0}\big(\rho_{1}(\cdots\rho_{N-1}(X)\cdots)\big).$
Nested risk measures are useful because they ensure time consistency. However, nested risk measures are typically hard to interpret. We note that typically in the RL literature, the terms ‘dynamic risk measure’ and ‘nested risk measure’ are used interchangeably; however, formally speaking, dynamic risk measures need not be time-consistent, nor have the nested structure.
Markov Risk Measures (Definition A.7): A Markov risk measure is a conditional, possibly-dynamic risk measure that is only conditioned on the information available at the most recent time step. Markov risk measures are useful because they enforce a one-step time dependence structure that makes them compatible with MDP-based RL solution methods.
@@ 3.2.1 Non-Continual Risk-Aware Reinforcement Learning
Traditionally, non-continual risk-aware RL works have aimed to optimize either a static, possibly-coherent risk measure (e.g. Mead et al. [2025]), or a nested, Markov, possibly-coherent risk measure (e.g. Ruszczyński [2010]). The primary appeal of optimizing static risk measures is that they are interpretable; however, they lack time consistency. Conversely, nested Markov risk measures are typically characterized as being difficult to interpret, but ensure time consistency. We note that there is no consensus as to which one of these two approaches is preferred; the trade-off between interpretability and time consistency reflects an open design choice in non-continual risk-aware RL. For conciseness, we will refer to the two aforementioned approaches in non-continual risk-aware RL as the ‘static’ and ‘nested’ approaches for the remainder of this text.

### v2

@@ Abstract
Continual reinforcement learning (continual RL) seeks to formalize the notions of lifelong learning and endless adaptation in RL. In particular, the aim of continual RL is to develop RL agents that can maintain a careful balance between retaining useful information and adapting to new situations. To date, continual RL has been explored almost exclusively through the lens of risk-neutral decision-making, in which the agent aims to optimize the expected long-run performance. In this work, we present the first formal theoretical treatment of continual RL through the lens of risk-aware decision-making, in which the behaviour of the agent is directed towards optimizing a measure of long-run performance beyond the mean. In particular, we show that the classical theory of risk measures, widely used as a theoretical foundation in non-continual risk-aware RL, is, in its current form, incompatible with continual learning. Then, building on this insight, we extend risk measure theory into the continual setting by introducing a new class of ergodic risk measures that are compatible with continual learning. Finally, we provide a case study of risk-aware continual learning, along with empirical results, which show the intuitive appeal of ergodic risk measures in continual settings.
@@ 1 Introduction
To date, the notion of continual RL has been studied almost exclusively under a risk-neutral lens, such that the behaviour of the agent is directed towards optimizing some expectation-based measure of long-run performance (e.g. the average per-step reward or the expected return). In this work, we present the first formal theoretical exploration of continual RL through the lens of risk-aware decision-making, in which the behaviour of the agent is directed towards optimizing a measure of long-run performance beyond the mean. Through this exploration, we seek to provide a theoretically-grounded formalism through which we can frame the notion of risk-awareness in a continual setting. Namely, the first half of our exploration is guided by the following questions:
  • $\triangleright$
    What does it mean for an agent to be risk-aware in a continual setting?
  • $\triangleright$
    What are the conditions needed to enable an agent to be risk-aware in a continual setting?
  • $\triangleright$
    What are the implications of risk-awareness as it relates to the stability-plasticity dilemma?
To answer these questions, we propose a risk-aware generalization of the continual RL framework proposed in Abel et al. [2025], where we formalize risk-aware continual RL as an ongoing exchange of actions and observations between an agent and an environment, such that the agent receives observations from the environment, processes those observations into risk-aware observations based on its risk attitude (or risk tolerance), then emits actions back to the environment based on the risk-aware observations. We then leverage this risk-aware framework to propose a set of axioms that establish the conditions needed to enable an agent to be risk-aware in a continual setting.
Figure 1: A comparison of: a) the continual RL framework proposed in Abel et al. [2025], and b) the risk-aware generalization of the framework proposed in this work.
The second half of our exploration is then aimed at examining existing frameworks and methodologies used for non-continual risk-aware RL to see how compatible they are with our proposed framework. To this end, we first examine the classical theory of risk measures (e.g. see Chapter 6 of Shapiro et al. [2009]), which has served as a theoretical foundation for risk-aware decision-making in non-continual RL, and show that, in its current form, it is incompatible with continual learning. Then, building on this insight, we extend risk measure theory into the continual setting by proposing a new class of ergodic risk measures that are compatible with continual learning.
Finally, using the well-known average-reward Markov decision process (MDP) formulation [Puterman, 1994] as a basis, we provide a case study, along with numerical results, which show the intuitive appeal of ergodic risk measures in continual settings. Altogether, these contributions provide, to the best of our knowledge, the first formal theoretical treatment of risk-aware decision-making in a continual (i.e., lifelong) learning setting.
@@ 2.1 Continual Reinforcement Learning
The notions of lifelong learning and endless adaptation in the context of RL have long been studied under different names and perspectives. In recent years, several works (e.g. Khetarpal et al. [2022], Abel et al. [2023], Kumar et al. [2025]) have attempted to unify and frame these diverse sets of works as instances of continual RL. Some of the more common types of RL-related works that can be interpreted as being instances of continual RL include the study of the loss of plasticity in deep RL agents (e.g. Abbas et al. [2023], Dohare et al. [2024]), transfer learning (e.g. Abel et al. [2018], Gimelfarb et al. [2021]), and decision-making in non-stationary environments (e.g. Dick et al. [2014], Luketina et al. [2022]). In this work, we place a great emphasis on the notion of plasticity, which has been studied extensively in prior works such as Raghavan and Balaprakash [2021], Chen et al. [2023], and Abel et al. [2025]. In our proposed framework, we build on the definition of plasticity (and subsequent continual RL framework) proposed in Abel et al. [2025].
@@ 2.2 Risk-Aware Reinforcement Learning
The notion of risk-aware learning and decision-making in the context of RL has been studied under various theoretical frameworks, from the well-established expected utility framework [Howard and Matheson, 1972], to the more contemporary framework of risk measures (e.g. Chapter 6 of Shapiro et al. [2009]). In this work, we focus on the latter framework, which originated in the finance literature (e.g. Rockafellar and Uryasev [2000]), but has since been widely integrated into RL-based works (e.g. Bäuerle and Ott [2011]). Traditionally, non-continual risk-aware RL works have aimed to optimize either a static risk measure (e.g. Mead et al. [2025]), or a nested (dynamic) risk measure (e.g. Wang and Delage [2025]). The case study presented in this work primarily focuses on the conditional value-at-risk (CVaR) risk measure [Rockafellar and Uryasev, 2000], which has been studied extensively in the discounted setting (e.g. Bäuerle and Ott [2011], Mead et al. [2025]), and, to a lesser extent, in the average-reward setting (e.g. Xia et al. [2023], Rojas and Lee [2025]).
@@ 3.1 Continual Reinforcement Learning
Continual reinforcement learning (continual RL) can be viewed as an agent-environment interaction occurring over an infinite time horizon. In particular, following the framework proposed in Abel et al. [2025], this interaction can be modeled as an exchange of signals between an agent and an environment, where, at each discrete time step, $t$, the agent emits an action from an action-space, $A_{t}\in\mathcal{A}$, and the environment emits an observation from an observation-space, $O_{t}\in\mathcal{O}$ (see Figure 1a)). Accordingly, for a given agent and environment, we can define their interaction as a pair of sequences of random variables, $O_{a:b}\doteq(O_{a},O_{a+1},\ldots,O_{b-1},O_{b})$ and $A_{c:d}\doteq(A_{c},A_{c+1},\ldots,A_{d-1},A_{d})$, where $a,b,c,$ and $d$ denote discrete time indices.
One of the defining aspects of continual RL is the stability-plasticity dilemma, through which an agent must carefully balance the degree to which newly acquired information affects its behaviour relative to previously learned knowledge. In particular, the plasticity, $\mathfrak{P}$, of a given agent relative to a given environment with respect to the time intervals $a:b]\doteq[a,a+1,\ldots,b-1,b]$, $1\leq a\leq b$, and $c:d]\doteq[c,c+1,\ldots,d-1,d]$, $1\leq c\leq d$, can be defined as follows [Abel et al., 2025]:
$\mathfrak{P}_{\begin{subarray}{c}a:b\\ c:d\end{subarray}}\doteq\mathbb{I}(O_{a:b}\to A_{c:d}),$ (1)
where $\mathbb{I}(X_{a:b}\to Y_{c:d})\doteq\sum_{i=\max(a,c)}^{d}\mathbb{I}(X_{a:\min(b,i)};Y_{i}\mid X_{1:a-1},Y_{1:i-1})$ denotes the generalized directed information [Abel et al., 2025] between two sequences of random variables, $X_{a:b}$ and $Y_{c:d}$, and $\mathbb{I}(X;Y)$ denotes the mutual information between random variables $X$ and $Y$. In words, the above definition interprets plasticity as a measure of how much a sequence of observations from the environment influences a sequence of the agent’s actions. Accordingly, the stability-plasticity dilemma can be viewed as the dilemma associated with determining the right amount of plasticity.
@@ 3.2 Risk Measures
Let $(\Omega,\mathcal{F},\mathbb{P})$ denote a probability space, and let $\mathcal{X}$ denote a space of random variables of the form $X:\Omega\to\mathbb{R}$. A risk measure (e.g. Chapter 6 of Shapiro et al. [2009]) is a functional, $\rho:\mathcal{X}\times\mathcal{T}\to\mathbb{R}$, that assigns to each random variable, $X\in\mathcal{X}$, a real value representing the degree of risk associated with $X$ with respect to some risk attitude or risk tolerance, $\tau\in\mathcal{T}$, where $\mathcal{T}$ denotes a set that captures the agent’s risk attitude. In other words, a risk measure is a mapping that quantifies the risk associated with a random variable based on a given risk attitude. The precise interpretation of what ‘risk’ means depends on the risk measure used; different risk measures capture different aspects of a random variable. We note that it is commonplace to write $\rho(X)$ or $\rho_{\tau}(X)$ as shorthand for $\rho(X,\tau)$.
@@
In essence, one can think of a risk measure as any functional that captures distributional characteristics of a random variable, typically beyond just its mean. However, an emphasis is usually placed on deriving risk measures that satisfy certain mathematical properties that can be meaningful in risk-based decision-making contexts. In Appendix A, we provide formal definitions of the various classes of risk measures used in the context of RL, where each (non-mutually-exclusive) class of risk measures can be thought of as satisfying a specific set of mathematical properties.
@@ 3.2.1 Non-Continual Risk-Aware Reinforcement Learning
Traditionally, non-continual risk-aware RL works have aimed to optimize either a static risk measure (e.g. Mead et al. [2025]), or a nested risk measure (e.g. Wang and Delage [2025]). We now provide informal definitions of these two classes of risk measures, as well as an informal definition pertaining to the notion time consistency as it relates to risk measures (see Appendix A for formal definitions):
Static Risk Measures (Definition A.2): A static risk measure, $\rho_{N}$, evaluates risk at a fixed point in time, $N$. In RL-based contexts, static risk measures can be useful for quantifying the risk of the return at the end of an episode. The primary appeal of static risk measures is their interpretability.
Nested Risk Measures (Definition A.6): A nested risk measure is a type of dynamic risk measure (i.e., a sequence of one-step risk measures, $\{\rho_{t}\}_{t=0}^{N}$; see Definition A.4), that is constructed recursively from the one-step risk measures, such that the risk at time $n$, $\rho_{n}$, can be computed as $\rho_{n}(X)\doteq\rho_{n}\big(\rho_{n+1}(\cdots\rho_{N}(X)\cdots)\big)$. Nested risk measures are useful because they ensure time consistency (described below). In RL-based contexts, nested risk measures are also useful because they can induce Bellman-like recursions with appealing dynamic programming-like properties. However, nested risk measures are typically hard to interpret. We note that typically in the RL literature, the terms ‘dynamic risk measure’ and ‘nested risk measure’ are used interchangeably; however, formally speaking, dynamic risk measures need not be time-consistent, nor have the nested structure.
Time-Consistent Risk Measures (Definition A.5): A dynamic risk measure, $\{\rho_{t}(X)\}_{t=0}^{N}$, is said to be time-consistent if, for all $X,X^{\prime}\in\mathcal{X}$ and all $t<N$, we have that $\rho_{t+1}(X)\leq\rho_{t+1}(X^{\prime})\implies\rho_{t}(X)\leq\rho_{t}(X^{\prime})$. Time-consistent risk measures can be appealing because they ensure that if one future outcome is deemed less risky than another at some time step, $t$, then that same outcome is not deemed more risky than the other at any other time step. One way to think about time consistency in RL-based contexts is to ask the question: can the agent change its mind about how risky something is based on new information? If the answer is yes, then there exists a lack of time consistency.